Grade 12Mechanics

Gravitational potential at the centre of a uniform solid sphere of mass ‘M’ and radius ‘R’ is (if potential due to the sphere at the infinity is GM/R .)

Utkarsh mishra
8 Years agoGrade 12

1 Answer

Eshan
8 Years ago
Dear student,

Work done in bringing a massmfrom infinity to center of solid sphere is

\int_{\infty}^{R}(-\dfrac{GMm}{r^2})dr+\int_R^0(\dfrac{G(M(\dfrac{r^3}{R^2})m)}{r^2})dr
=-\dfrac{3}{2}\dfrac{GMm}{R}
Hence the potential at the center of the sphere=-\dfrac{3}{2}\dfrac{GM}{R}+V_{\infty}=-\dfrac{1}{2}\dfrac{GM}{R}